Brizendine, Inc., manufactures and sells two products: Product R0 and Product B6. The production of Product R0 is 900 units and of Product B6 is 700 units. The company has an activity-based costing system with the following activity cost pools, activity measures, and expected activity:

Estimated Expected Activity
Activity Cost Pools Activity Measures Overhead Cost Product R0 Product B6 Total
Labor-related DLHs $ 507,394 7,200 3,500 10,700
Product testing tests 9,464 300 500 800
Order size MHs 501,205 3,100 2,800 5,900
$ 1,018,063

The overhead applied to each unit of Product R0 under activity-based costing is closest to: (Round your intermediate calculations to 2 decimal places.)

Multiple Choice

  • $675.91 per unit

  • $292.61 per unit

  • $761.20 per unit

  • $636.29 per unit
Answer #1

Total overhead allocated to RO = (507394*7200/10700)+ (9464*300/800)+ (501205*3100/5900)

= $608318

Overhead applied to each unit of product RO = 608318/900 = $675.91 per unit